Math 120 A, B - Winter 2011 Mid-Term Exam Number Two February 24 , 2011 Answers

There were two versions of the exam. Version A - Problem 1 involved 1000 meters of fencing.

1. (a) 67.6788 m (b) The radius of the circle should be zero.

2. (a) 0.06652 m (b) 5.0555^{∘}

3. (a) There are infinitely many such functions. One is f(x) = \frac{x + 4}{x − 2} (b) f^{−1}(x) = 2 − \sqrt{x − 3}

4. 447.84 years after 2005

Version B - Problem 1 involved 800 meters of fencing.

1. (a) 52.7311 m (b) The radius of the circle should be zero.

2. (a) 0.11104 m (b) 16.8517^{∘}

3. (a) There are infinitely many such functions. One is f(x) = \frac{x + 10}{x + 4} (b) f^{−1}(x) = 5 + \sqrt{7 + x}

4. 24.87 years after 1995